e^(x2+5x+6)=1
The use of the "e" value causes me confusion with these questions.

Answer by Edwin McCravy(20060)


ex2+5x+6 = 1
e is merely the base of natural logarithms. You don't even need
to know its value which is approximately 2.718.
Since you know that any number to the 0 power is 1
rewrite the 1 on the right side as e0
ex²+5x+6 = e0
Now you can simply set the exponents of e equal to each other:
x² + 5x + 6 = 0
And you can easily solve that by factoring.
Answers: -2 and -3
Edwin